# Stimulating the Diffusion Model for Image Denoising via Adaptive Embedding and Ensembling

### v3

@@ 5.2 Gaussian Image Denoising
${}_{+}$ MIRNet MPRNet DeamNet NAFNet Uformer Restormer
@@
${}_{+}$ MIRNet MPRNet DeamNet NAFNet Uformer Restormer
@@
In this section, we conduct Gaussian image denoising experiments on synthetic benchmark datasets. Image denoising is commonly evaluated using the distortion-based PSNR metric with noise levels $\sigma=15,25,50$. However, our method is robust to much higher noise levels, excelling in both distortion-based and perception-based metrics. Hence, we perform classical Gaussian denoising experiments as well as robust Gaussian denoising experiments.
Subsequently, we will provide a detailed explanation of our comparative methods. DnCNN [7] and DRUNet [6] are considered classical methods, while Restormer [5] and ART [71] represent the current state-of-the-art approaches. For DnCNN [7] and DRUNet [6], we retrain them following their specified training details. However, it’s worth noting that retraining Restormer [5] and ART [71] is a resource-intensive process. Their publicly available pre-trained models are not designed to handle noise levels with standard deviations exceeding 50. Therefore, following the approach of [6], we multiply by a constant to ensure that the standard deviation is 50.
@@ 5.3 Real-world Image Denoising
We compare our method with both supervised and unsupervised methods. As our method is essentially an unsupervised solution for real-world denoising, which doesn’t require training on pairs of noisy-clean real-world images. Secifically, we evaluate different methods on three datasets: CC [72], PolyU [73], and FMDD [74], following [53, 81, 32, 80]. These datasets contain 16 paired sRGB images, 100 paired sRGB images, and 48 paired raw images, respectively.
@@
Subsequently, we will provide a detailed explanation of the comparative methods. For supervised methods, ${}_{+}$ [75] is the only and the latest generative method that can be directly employed for denoising. In addition, all the supervised methods are pre-trained on real-world dataset SIDD [84] and borrowed from their officially released versions. For unsupervised methods, the results of AP-BSN [30], LG-BPN [31], ZS-N2N [80], R2R [32] are reproduced and evaluated by ourselves, since these methods are not evaluated on all the datasets and metrics we use in their original paper. The results of N2V [28], N2S [29] and S2S [79] are reported from R2R [32] and ScoreDVI [81]. Since we reproduce more advanced methods AP-BSN [30] and LG-BPN [31] for N2V [28] and N2S [29]. In addition, S2S [79] requires too many iterations ($4.5\times 10^{5}$) to produce a single image which takes around $4\sim 5$ hours to process an image of size $512\times 512$. For our method, the noise level of the latent noisy image after noise transformation is pre-defined as a constant for CC [72], PolyU [73], and FMDD [74], following NN method [53]. In practice, fine-tuning the noise level parameter could potentially lead to even better performance for our method.
@@
Table III and Figure 7 demonstrate our significant advantages. The table III segregates supervised methods in the top row and unsupervised methods in the bottom row. In contrast to supervised methods, our unsupervised approach yields significantly higher PSNR and SSIM scores. Supervised methods, trained on the SIDD dataset [84], suffer from poor generalization, rendering them highly vulnerable to out-of-distribution data. As a result, their outcomes consistently exhibit significant residual noise. In contrast to unsupervised methods, our results demonstrate a substantial improvement. AP-BSN [30] and LG-BPN [31] often yield overly blurred outputs, accompanied by texture distortion, whereas our results maintain remarkably clear textures. R2R [32] and ZS-N2N [80] appear to grapple with complete noise removal, leaving residual noise in the final results. In contrast to the generative method, our results surpass ${}_{+}$ [75] comprehensively across both metrics and visual performance. These experiments unequivocally underscore the adeptness of our method in handling denoising tasks, achieving high perceptual quality while introducing minimal distortion.
@@
The abnormal situation observed in the FMDD [74] dataset concerning the LPIPS metric is worth discussing. It is notable that our method and ${}_{+}$ [75], both generative methods, perform well on LPIPS for CC [72] and PolyU [73] but fail for FMDD [74]. Additionally, our method and LG-BPN [31] perform well on PSNR and SSIM for FMDD [74] but not on LPIPS. This situation appears abnormal. We hypothesize that it may be due to the perception-distortion curve phenomenon [19]. The perception-distortion curve implies a contradiction between distortion-based metrics and perception-based metrics. Therefore, outstanding performance on distortion-based metrics can not align with outstanding perception-based metrics. This explanation is consistent with our results on CC [72], where our DMID-d does not achieve the second-best results on LPIPS. Despite the presence of such an abnormal situation, our method still stands out among unsupervised methods. Disregarding this abnormal situation, our method shines brightly even when compared to various supervised methods.
@@
TABLE IV: Ablation study of Embedding method. Setting1 is referred to whether perform noise transformation and Setting2 is referred to whether convert the image to the intermediate state $x_{N}$ of diffusion model.
CC [72] FMDD [74]
Setting1 Setting2 PSNR / SSIM / LPIPS PSNR / SSIM / LPIPS
@@ 5.4.1 For embedding method
Without converting the image into the intermediate state $x_{N}$ of the diffusion model, our method cannot function. Therefore, we consistently convert the image into the intermediate state $x_{N}$, whether performing noise transformation or not. The results are presented in Table IV, and it is evident that noise transformation brings about significant improvement.
This experiment demonstrates the significance of noise transformation for our method. With noise transformation, our approach becomes versatile in addressing various noise types. (b). Influence of $N$. We conduct experiments on McMaster with noise level $\sigma=25$ using different choices of $N$ with a fixed number of sampling times $S_{t}=1$ and repetition times $R_{t}=1$.
@@
Since LPIPS is more tolerant to noise than to blurring [83], the highest LPIPS score is frequently achieved when denoising is slightly incomplete. When $\eta\approx 1$ at timesetp $N$, the timesetp $N$ is denoted as $N_{\text{optimal}}$. As for our method, the optimal LPIPS is generally obtained within the range of $N_{\text{optimal}}-1$ to $N_{\text{optimal}}-5$ as shown in Figure 9. In our experiments, we just set $N$ to be $N_{\text{optimal}}$, and further tuning could potentially yield even better LPIPS results. This phenomenon, in which the optimal LPIPS score is not attained at $N_{\text{optimal}}$ while PSNR is achieved at $N_{\text{optimal}}$, further validates the credibility of our approach. It underscores the accuracy of our $N$ calculation.
Furthermore, our method demonstrates robustness, showing favorable denoising outcomes when $\eta$ varies within the range of $1\pm 10\%$ as shown in Figure 9. This experiment demonstrates the necessity of embedding into intermediate states and the appropriate choice of $N$ for our method. Without embedding into intermediate states, our method cannot perform denoising. With the appropriate choice of $N$, our denoising performance reaches its maximum potential.
@@
5.4.2 For ensembling method
Our ensembling method adaptively adjusts distortion and perception based on requirements by adjusting the sampling times and repeating the inference process. Therefore, for the ensembling method, we separately evaluate the impact of different values for sampling times $S_{t}$ and repetition times $R_{t}$.
@@ 5.4.2 For ensembling method
$\sigma=50$ 28.76 / 0.8950 / 0.163 29.20 / 0.9030 / 0.158 29.92 / 0.9155 / 0.121
$\sigma=100$ 26.12 / 0.8395 / 0.261 26.40 / 0.8484 / 0.249 27.01 / 0.8621 / 0.211
$\sigma=150$ 24.70 / 0.8036 / 0.316 24.77 / 0.8085 / 0.308 25.39 / 0.8233 / 0.274
$\sigma=200$ 23.66 / 0.7737 / 0.362 23.61 / 0.7763 / 0.352 24.28 / 0.7931 / 0.305
$\sigma=250$ 22.81 / 0.7465 / 0.405 22.71 / 0.7487 / 0.387 23.44 / 0.7666 / 0.360
5.5 Comparative Analysis and Extension to Diffusion-Based Methods
@@
Firstly, we conduct experiments compared with other diffusion-based image restoration methods. DDRM [48] and DDNM [49] are the representative diffusion-based image restoration methods. Tables VII and VI showcase the results of Gaussian denoising on ImageNet and real-world denoising on CC [72], PolyU [73], and FMDD [74]. As DDRM [48] and DDNM [49] do not focus on image denoising, we do not compare them in the previous Section 5.2 and Section 5.3, and the results in the table are produced by ourselves. The hyperparameters of noise level used in DDRM [48] and DDNM [49] are estimated using the method [85]. Other hyperparameters are employed their recommended settings. Due to their poor performance on real-world denoising, we upgrade DDRM and DDNM to “DDRM+Clean” and “DDNM+Clean”, respectively. Here, for “DDRM+Clean” and “DDNM+Clean”, we introduce the clean image as auxiliary information to calculate the noise level for the corresponding noisy image. However, their performance is still limited.
As depicted in Figure 13, both DDRM [48] and DDNM [49] struggle to handle image denoising tasks. That is because these methods are not well-suited for image denoising tasks. As discussed in Section 2.2, they can not circumvent the input inconsistency and overlook the content inconsistency. In contrast, our method is designed to tackle the two problems. Firstly, our method involves embedding the noisy image with different noise into an intermediate state, from which we subsequently sample. This intermediate state is closely tied to the noise level, and our embedding technique allows for the denoising of a wide array of noise types. Additionally, we establish a clear and noise-dependent starting point for the sampling process. Secondly, our method introduces an innovative ensembling method. This ensembling method enables optimization based on either distortion-based or perception-based requirements. Notably, we provide a comprehensive comprehension of how both distortion-based and perception-based quality are influenced, and we present strategies for fine-tuning these aspects according to specific requirements. In addition, our DMID method represents a general strategy for denoising within diffusion models, without imposing restrictions on the sampling strategy, as explained in Section 4.1. Thus, our method can extend to other strategies and help them avoid input inconsistency and content inconsistency.
@@
The visual results are presented in Figure 14. The versions featuring our embedding method and ensembling method (referred to ”DM +Ours”), outperform their original counterparts DDRM [48] and DDNM [49] easily. Additionally, ”DM +Ours” employs significantly fewer sampling times compared to DDRM [48] and DDNM [49], respectively. For example, the sampling times $S_{t}$ are 3 for ”DM +Ours”, whereas DDRM [48] and DDNM [49] require 20 and 100 sampling times, respectively, on the PolyU [73] dataset. This difference arises because DDRM [48] and DDNM [49] start sampling from pure Gaussian noise for image denoising. The generated intermediate image $x_{N}$ with a noise level similar to that of the original input noisy image $y$ does not align with the content and information of the original input noisy image $y$ for DDRM [48] and DDNM [49]. Consequently, their results not only exhibit poor performance but also demand a larger number of sampling times. This experiment further emphasizes the superiority, effectiveness, and scalability of our method.

### v5

@@ 5.2 Gaussian Image Denoising
DANet+ MIRNet MPRNet DeamNet NAFNet Uformer Restormer
@@
DANet+ MIRNet MPRNet DeamNet NAFNet Uformer Restormer
@@ 5.3 Real-world Image Denoising
We compare our method with both supervised and unsupervised methods. As our method is essentially an unsupervised solution for real-world denoising, which does not require training on pairs of noisy-clean real-world images. Secifically, we evaluate different methods on three datasets: CC [79], PolyU [80], and FMDD [81], following [57, 88, 37, 87].
@@
Subsequently, we will provide a detailed explanation of the comparative methods. For supervised methods, DANet+ [82] is the only and the latest generative method that can be directly employed for denoising. In addition, all the supervised methods are pre-trained on real-world dataset SIDD [89] and borrowed from their officially released versions. For unsupervised methods, the results of AP-BSN [35], LG-BPN [36], ZS-N2N [87], R2R [37] are reproduced and evaluated by ourselves, since these methods are not evaluated on all the datasets and metrics we use in their original paper. The results of N2V [33], N2S [34] and S2S [86] are reported from R2R [37] and ScoreDVI [88].
@@
Table III and Figure 7 demonstrate our significant advantages. The Table III segregates supervised methods in the top row and unsupervised methods in the bottom row. In contrast to supervised methods, our unsupervised approach yields significantly higher PSNR and SSIM scores. Supervised methods, trained on the SIDD dataset [89], suffer from poor generalization, rendering them highly vulnerable to out-of-distribution data. As a result, their outcomes consistently exhibit significant residual noise. In contrast to unsupervised methods, our results demonstrate a substantial improvement. AP-BSN [35] and LG-BPN [36] often yield overly blurred outputs, accompanied by texture distortion, whereas our results maintain remarkably clear textures. R2R [37] and ZS-N2N [87] appear to grapple with complete noise removal, leaving residual noise in the final results. These experiments unequivocally underscore the adeptness of our method in handling denoising tasks, achieving high perceptual quality while introducing minimal distortion.
@@
The abnormal situation observed in the FMDD [81] dataset concerning the LPIPS metric is worth discussing. It is notable that our method and DANet+ [82], both generative methods, perform well on LPIPS for CC [79] and PolyU [80] but fail for FMDD [81]. Additionally, our method and LG-BPN [36] perform well on PSNR and SSIM for FMDD [81] but not on LPIPS. This situation appears abnormal. We hypothesize that it may be due to the perception-distortion curve phenomenon [24]. The perception-distortion curve implies a contradiction between distortion-based metrics and perception-based metrics. Therefore, outstanding performance on distortion-based metrics can not align with outstanding perception-based metrics. This explanation is consistent with our results on CC [79], where our DMID-d does not achieve the second-best results on LPIPS. Despite the presence of such an abnormal situation, our method still stands out among unsupervised methods. Disregarding this abnormal situation, our method shines brightly even when compared to various supervised methods.
@@
TABLE IV: Results of the ablation study on the Embedding method with whether perform noise transformation.
CC FMDD
Noise Transformation PSNR / SSIM / LPIPS PSNR / SSIM / LPIPS
@@ 5.4.1 For embedding method
Without converting the image into the intermediate state $x_{N}$ of the diffusion model, our method cannot function. Therefore, we consistently convert the image into the intermediate state $x_{N}$, whether performing noise transformation or not. The results are presented in Table IV, and it is evident that noise transformation brings about significant improvement. This experiment demonstrates the importance of noise transformation for our method.
(b). Influence of $N$. We conduct experiments on McMaster with noise level $\sigma=25$ using different choices of $N$ with a fixed number of sampling times $S_{t}=1$ and repetition times $R_{t}=1$.
@@
Since LPIPS is more tolerant to noise than to blurring [78], the highest LPIPS score is frequently achieved when denoising is slightly incomplete. When $\eta\approx 1$ at timesetp $N$, this optimal timesetp $N$ is denoted as $N_{o}$. As for our method, the optimal LPIPS is generally obtained within the range of $N_{o}\!\!-\!1$ to $N_{o}\!\!-\!5$ as shown in Figure 9. In our experiments, we just set $N$ to be $N_{o}$, and further tuning could potentially yield even better LPIPS results. This phenomenon, in which the optimal LPIPS score is not attained at $N_{o}$ while PSNR is achieved at $N_{o}$, further validates the credibility of our approach. It underscores the accuracy of our $N$ calculation. Furthermore, our method demonstrates robustness, showing favorable denoising outcomes when $\eta$ varies within the range of $1\pm 10\%$ as shown in Figure 9.
This experiment demonstrates the necessity of embedding into intermediate states and the appropriate choice of $N$ for our method. Without embedding into intermediate states, our method cannot perform denoising. With the appropriate choice of $N$, our denoising performance reaches its maximum potential.
5.4.2 For ensembling method
Our ensembling method adjusts distortion and perception based on requirements by adjusting the sampling times and repeating the inference process. Therefore, for the ensembling method, we separately evaluate the impact of different values for sampling times $S_{t}$ and repetition times $R_{t}$.
@@ 5.4.2 For ensembling method
5.5 Comparative Analysis and Extension to Diffusion-Based Methods
@@
$\sigma=50$ 28.76 / 0.8950 / 0.163 29.20 / 0.9030 / 0.158 29.90 / 0.9157 / 0.114
$\sigma=100$ 26.12 / 0.8395 / 0.261 26.40 / 0.8484 / 0.249 27.00 / 0.8626 / 0.201
$\sigma=150$ 24.70 / 0.8036 / 0.316 24.77 / 0.8085 / 0.308 25.39 / 0.8236 / 0.263
$\sigma=200$ 23.66 / 0.7737 / 0.362 23.61 / 0.7763 / 0.352 24.25 / 0.7915 / 0.295
$\sigma=250$ 22.81 / 0.7465 / 0.405 22.71 / 0.7487 / 0.387 23.44 / 0.7675 / 0.346
Figure 13: The comparison of runtime on the ImageNet dataset with noise level $\sigma=250$. Notably, our method consistently yields optimal results under the same runtime.
@@
Firstly, we conduct experiments compared with other diffusion-based image restoration methods. DDRM [50] and DDNM [51] are the representative diffusion-based image restoration methods. Tables VII and VI showcase the results of Gaussian denoising on ImageNet and real-world denoising on CC [79], PolyU [80], and FMDD [81]. Our method consistently yields optimal results under the same runtime as shown in Figure 13. As DDRM [50] and DDNM [51] do not focus on image denoising, we do not compare them in the previous Section 5.2 and Section 5.3, and the results in the table are produced by ourselves. The hyperparameters of noise level used in DDRM [50] and DDNM [51] are estimated using the method [90]. Other hyperparameters are employed their recommended settings. Due to their poor performance on real-world denoising, we introduce the clean image as auxiliary information to calculate the noise level and upgrade DDRM and DDNM to “DDRM+Clean” and “DDNM+Clean”, respectively. However, the performance is still limited.
As depicted in Figure 14, both DDRM [50] and DDNM [51] struggle to handle image denoising tasks. That is because these methods are not well-suited for image denoising tasks. As discussed in Section 2.2, they can not circumvent the input inconsistency and overlook the content inconsistency.
@@
The visual results are presented in Figure 15. The versions featuring our embedding method and ensembling method (referred to ”DM +Ours”), outperform their original counterparts DDRM [50] and DDNM [51] easily. Additionally, ”DM +Ours” employs significantly fewer sampling times compared to DDRM [50] and DDNM [51], respectively. For example, the sampling times $S_{t}$ are 3 for ”DM +Ours”, whereas DDRM [50] and DDNM [51] require 20 and 100 sampling times, respectively, on the PolyU [80] dataset. DDRM and DDNM require more runtime because DDRM [50] and DDNM [51] start sampling from standard Gaussian noise for image denoising. Moreover, when DDRM and DDNM produce the intermediate image $x_{N}$ with a noise level comparable to that of the original input noisy image $y$, the content and information of $x_{N}$ do not align with those of the original input noisy image $y$. Consequently, DDRM [50] and DDNM [51] not only exhibit poor performance but also demand a larger number of sampling times. This experiment further emphasizes the superiority, effectiveness, and scalability of our method.
6 Discussion
In this paper, we introduce a method that utilizes the diffusion model pre-trained for image synthesis. At present, there is no established paradigm for effectively leveraging the diffusion model for image denoising. The comparison between trainable and pre-trained solutions raises an intriguing question that has yet to be thoroughly explored. In the subsequent paragraphs, we will analyze the merits and drawbacks of both approaches and aim to draw a conclusion.
The pre-trained solution employs the same pre-trained diffusion model for universal image denoising. Utilizing pre-trained diffusion models eliminates the need for training and provides extensive image priors, enhancing denoising capabilities. Despite these advantages, challenges related to input inconsistency and content inconsistency are required to be corrected.
The trainable solution customizes diffusion models for various noise types and datasets [52], improving the handling of specific noise. However, complexities arise from the unknown distribution of real noise and the sensor-specific nature of noise [91], leading to improbability in customizing specific models for real noise and inefficiencies in training different models for various noise. In addition, diffusion models are delicate and only excel with Gaussian noise [92, 93]. Customizing diffusion models and altering the noise distribution leads to suboptimal performance [92, 93].
In summary, our solution which employs pre-trained diffusion models offers a flexible solution for handling various types of noise. However, some modifications are required to adapt the pre-trained diffusion model. On the other hand, trainable diffusion models provide a solution for customizing the training of different diffusion models based on distinct noise profiles. Nevertheless, challenges arise in customizing the training process and optimizing the denoising capabilities.